# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ VI Generation of magic States
Magic states are an important ingredient of universal FTQC as they allow implementation of non-Clifford operations while themselves remaining distillable with only Clifford gates Bravyi and Kitaev (2005). A popular gate set which generates the Clifford group is {$CNOT$, $H$, $S$}, where $CNOT$ is the controlled NOT gate; $H$ is the Hadamard gate, and $S$ is the phase gate. The non-Clifford T gate usually completes the universal gate set Bravyi and Kitaev (2005). Provided with a source of physical magic states, one can distill a resource state that enables teleportation of non-Clifford operations into the QEC substrate at arbitrarily low error rates. Improving magic state quality by increasing the fidelity with respect to any of the |$\ket{H}$ or |$\ket{T}$ states reduces the physical resources required for a given logical error rate in the quantum computation. A promising method to produce a GKP magic state uses GKP error correction on the vacuum state Baragiola et al. (2019), implying that GKP qubits and encoded Clifford operations are sufficient for universal QC Baragiola et al. (2019); Calcluth et al. (2022). However, simulations show that a reasonable success rate with existing methods requires the effective squeezing of GKP states to be extraordinarily high, approaching $20$ dB (see Fig. 8), to create distillable magic states in one error correction round. In other architectures, producing magic states with practical success rates require the GKP states to have very high effective squeezing Baragiola et al. (2019); Takase et al. (2024), well above the FT threshold. In this section we provide a method to circumvent this issue and produce magic states with GKP resources that have effective squeezing levels near the FT threshold.
@@
The distillation infidelity threshold for H-type states is approximately 14.7%. The plots in Fig. 8 show a deterministic simulation of CV-QEC on input vacuum or cat states with post-selection on quadrature measurement outcomes of zero. Note that this post-selection, done to ease simulations, will slightly underestimate the output squeezing of the bred GKP states Weigand and Terhal (2018). We see that when performing CV-QEC on the vacuum mode (purple data points) one obtains infidelities above the distillation threshold (of 14.7%) when using GKP states with squeezing of 13dB or less as the resource state for CV-QEC. On the other hand, with effective squeezing greater than 12dB and cat states with amplitude of $\sim 5.1$, magic states with infidelities below the distillation threshold can be obtained.
@@
In practice, the cat states produced by PhANTM have random amplitudes, and they must be squeezed in order to obtain the correct spacing for GKP breeding before magic state generation. This means that our GKP states have a random squeezing distribution. Further, as Fig. 8 shows, the optimal input cat amplitude for the error-corrected mode is between 4.9 and 5.5. Because our cats have random amplitudes and squeezing levels, only a few will hit this required target, although PhANTM can be engineered to emit cat states with mean squeezing levels and amplitudes centered on the requirements. Monte Carlo simulations were used to determine the success rate of producing a magic state when random cats from PhANTM were used to provide the resource states, both for GKP breeding and the cat state upon which CV QEC was performed. The simulations use Qutip Johansson et al. (2013) for density matrix calculations and Strawberry Fields Killoran et al. (2019) for homodyne detection simulation. Fig. 10 shows histograms of the infidelities obtained from each magic state generation attempt, where the fidelity is measured with respect to an ideal $H$-type magic state. The cat states were produced using Monte Carlo PhANTM simulations with a cluster state squeezing of 13.5 dB, 20 PhANTM steps, and 8 photon-subtraction attempts made before homodyne detection and teleportation to the next PhANTM time step. When using vacuum states as the error-corrected mode, we see a success rate of 0.003 after 1000 iterations (brown counts in the histogram). On the other hand, when using cat states as the error-corrected mode, we see a success rate of 0.048 after 1000 magic state generation attempts, an improvement of more than factor of 10, which reduces resources required to produce a magic state by a factor 10 relative to using vacuum alone. This simulation specifically targeted $H$-type states, which enable teleportation of $T$ gates within the architecture. These success rates mean that magic factories will not require the majority of qubit resources in surface code substrates using lattice surgery Litinski (2019a).
@@
VII Encoding Logical Qubits in a Surface Code
GKP encoding can protect against small displacement errors, but larger displacements ($>\frac{\pi}{2}$) will result in an error. Depending on the quadrature, this can be either a Pauli X error ($q$ quadrature) or a Pauli Z error ($p$ quadrature). These logical qubit errors require additional qubit-level error-correction. Many error correcting codes have been studied in the context of GKP qubits Bourassa et al. (2021); Stafford and Menicucci (2023); Raveendran et al. (2022); Zhang et al. (2021). We will focus on topological error correction based on a family of codes known as surface codes, which have been shown to have high thresholds ($\approx 1\%$ under circuit level noise Wang et al. (2011) or $\approx 10\mathrm{dB}$ with GKP states Bourassa et al. (2021)) and well-studied logical gates Litinski (2019b).
As mentioned earlier, photonic CVQC is well-suited to MBQC. Circuit model codes known as Calderbank-Steane-Short(CSS) codes, where X and Z errors are detected independently, can be translated into a cluster state for fault-tolerant MBQC via a process known as foliation Bolt et al. (2016), which loosely maps the data qubits in the circuit model code to linear cluster states. Surface codes are examples of CSS codes, and the resulting foliated 3D cluster state is often called the Raussendorf-Harrington-Goyal (RHG) lattice after the authors who first introduced it Raussendorf et al. (2007). The unit cell of this cluster state is shown in Fig. 11. The product of $X$-basis measurements (equivalently $p$-basis homodyne measurements for GKP qubits) on all face qubits is a stabilizer of the cluster state. In the error free case, measuring stabilizers will result in +1 eigenvalue and errors will result in $-1$ measurement results for any stabilizer that they anti-commute with. A second unit cell can be defined by shifting the base unit cell by 1/2 in all directions. Here, the edge qubits of the initial unit cell become face qubits of the second unit cell. These two lattices are known as primal and dual lattices, respectively. By measuring the stabilizers of both lattices, errors on all qubits can be detected. Logical operators in foliated cluster states are formed by correlation surfaces that map logical operators from the underlying code from one end of the cluster state to the other. To implement logical X and Z operations, correlation surfaces of the RHG lattice are given by 2D sheets of qubits which span the lattice and intersect primal and dual unit cells, respectively. The code distance is the shortest chain of errors that can result in a logical error, which is equivalent to the shortest error chain that can span the RHG lattice. The threshold theorem states that provided the physical error rate of the underlying qubits is below some threshold, we can arbitrarily suppress the logical error rate by increasing the code distance Knill et al. (1998).
@@ VII Encoding Logical Qubits in a Surface Code
The popular Clifford + T gate set maps naturally to the surface code. Clifford gates and teleportation based T gates can be implemented by lattice surgery Litinski (2019b). This approach defines logical qubits as patches, which are subsections of the overall surface code lattice, with logical operators identified on the edges. In MBQC, this corresponds to volumes of the RHG lattice with correlation surfaces on the boundaries. Patches can be merged by measuring stabilizer checks between edges of the patches, thus performing a joint logical measurement of the two patches. Clifford gates and the T gate can be implemented with joint measurements including ancilla qubits or magic states. In fact, in Ref. Litinski (2019b), it was shown how all Clifford gates can be commuted through the T gates and measurements, resulting in circuits that just consist of mulit-qubit joint measurements including a magic state. Lattice surgery operations can also be mapped to changing measurement bases (homodyne angles) in fault-tolerant MBQC Herr et al. (2018).
This gate set is well studied and optimized, however our architecture can be adapted to other gate sets Litinski (2019b); Gidney and Fowler (2019) and other error correcting codes Walshe et al. (2024). In this paper we aim to quantify Gaussian squeezing resources needed to demonstrate FT thresholds with the RHG lattice serving as the example code. We leave optimization of the error correcting code and gate set for future work.
VII.1 Macronode GKP lattice construction
@@
We can construct the lattice directly, where the nodes are GKP $\ket{+}$ states and the edges are CV $C_{Z}$ gates, which can be implemented deterministically with linear optics and squeezing. This approach has two main draw backs. Firstly, while it has been demonstrated previously Yoshikawa et al. (2008), in-line squeezing is generally difficult to realize experimentally. Second, ideal $C_{Z}$ gates are not physical, requiring infinite squeezing. Approximating the gate with finite squeezing will add noise to the GKP qubits. Alternatively, any lattice can be constructed with passive interferometers starting with GKP Bell pairs, where the number of physical modes per macronode is the valency of the underlying canonical graph state and the GKP Bell pairs are distributed to match the entanglement in the canonical graph state. Relating measurements on the physical modes with effective measurements on the underlying lattice requires a list of corrections known as a dictionary protocol Alexander et al. (2014), which we describe in Appendix XI.6. Note, that in the following we use the terms GKP state and sensor state interchangeably, always meaning either a sensor state or one half of a GKP Bell pair formed by interfering sensor states.
@@
We wish to design a modular construction of the required lattice. To this end, we identify a unit cell consisting of 24 sensor states or 12 GKP Bell pairs that can be tiled to construct a macronode RHG lattice of arbitrary size. From this unit cell, four GKP nodes are sent to neighboring unit cells, and correspondingly four are received. Two of the modes are delayed to form temporal links. This is illustrated in Fig. 11. As the primal and dual layers are created in a single time-step, we avoid the requirement for adaptive optics that have often been used to account for the different entanglement structure in primal and dual layers Bourassa et al. (2021); Larsen et al. (2021). To teleport non-Clifford operations onto logical qubits within the surface code, physical magic states are injected into the RHG lattice Litinski (2019a); Herr et al. (2018). The physical magic state is interfered with a vacuum state at a beamsplitter and is used in place of one GKP Bell pair in the RHG macronode.
VIII FTQC Threshold Simulation
@@
Several CV FTQC architectures based on GKP qubits and the surface code exist in the literature and have been characterised by their GKP effective squeezing threshold. Refs. Fukui et al. (2018) and Bourassa et al. (2021) used constructions based on $C_{Z}$ gates and found thresholds of $10$ dB and $10.5$ dB respectively, however these ignored the additional noise from finite squeezing in physical gates. In Ref. Larsen et al. (2021) found a threshold of $12.7$ dB while accounting for finite squeezing. Ref. Tzitrin et al. (2021) proposed a macronode-based architecture which did away with the need for $C_{Z}$ gates and found a threshold of $10.1$ dB. Refs. Bourassa et al. (2021); Larsen et al. (2021); Tzitrin et al. (2021) all require fast reconfigurable optics either in the computation or in the preparation of GKP states with a sufficiently high success probability.
@@
In the following section, we analyze the fault-tolerant performance of our architecture. In order to allow comparison with other architectures, we first use a phenomenological error model used widely in the literature Fukui et al. (2018, 2021); Larsen et al. (2021), which models realistic GKP states as ideal GKP states acted on by a Gaussian displacement channel. Formally, this is represented as Tzitrin et al. (2021)
@@
Which has the effect of replacing the delta function teeth of an ideal GKP state with Gaussian distributions with a standard deviation, $\Delta$,where $\Delta$ describes the same broadening of the GKP teeth as $\Delta_{D}$ described in Sec. II Bourassa et al. (2021). We can link this to the GKP effective squeezing in dB as $S_{\mathrm{dB}}=-\log_{10}\left(\Delta^{2}\right)$ Bourassa et al. (2021). Despite its simplicity, this error model can describe many common noise sources such as finite squeezing and photon loss Tzitrin et al. (2021); Bourassa et al. (2021); Larsen et al. (2021); Noh and Chamberland (2020). We then move on to quantify the fault-tolerance performance in terms of the input Gaussian resource for PhANTM - the squeezing level of the multi-mode cluster state. The input for these simulations are distributions of GKP squeezings obtained from simulations of the proposed GKP generation pipeline.
@@
In both cases, to determine the logical error rate we track how the GKP effective squeezing propagates through the optical circuit used to generate the macronode lattice and use the propagated noise variances to simulate noisy homodyne measurements. We apply the corresponding post-processing to reduce the physical macronode measurements to the effective noisy measurements on the canonical lattice, as shown in Appendix XI.6. With these noisy measurements we determine the probability of a qubit level error on each node of the canonical lattice, see Appendix XI.7 for more detail. These probabilities are used to sample a set of errors and generate the corresponding syndrome of the RHG lattice. This syndrome is then decoded using the pyMatching implementation of the minimum weight perfect matching decoder Higgott and Gidney (2023). This simulation method is described in more detail in Appendix XI.8 and is similar to the method used in Ref Bourassa et al. (2021), adapted to the macronode approach in Ref. Tzitrin et al. (2021) and for our specific GKP generation method.
VIII.1 Phenomenological error model
@@
Here, we present results which tailor the phenomenological noise model to the PhANTM and adaptive breeding approach to generate GKP states. As described in Sec. V, generating GKP states from PHANTM and breeding results in GKP states that are inherently asymmetric due to the different physical processes that determine the squeezing in the GKP quadratures. Note that this is equivalent to grid states with asymmetric spacing in the two quadratures Stafford and Menicucci (2023) which has also been studied in Walshe et al. (2024). This feature was incorporated into our simulation method by defining $p$ and $q$ noise variances separately for each GKP state. While we can also set each GKP state independently, for simplicity we assume that that the $p$ and $q$ noise are identical for each GKP state. To determine the 2D threshold curve for unbalanced quadrature noise we employ the method used in Ref. Bartolucci et al. (2021). By expressing the $p$ and $q$ noise in terms of a dummy variable, $\Delta_{p}=c_{p}\times x$, $\Delta_{q}=c_{q}\times x$, we fix the ratio of $p$ and $q$ noise via $c_{p}/c_{q}$ and use $x$ to find the threshold for a given ratio.
@@
Fig. 12 shows the resultant surface, where we have shaded the correctable region. The surface is symmetric with marginal squeezing thresholds of $7.2\mathrm{dB}$. With balanced squeezing in each quadrature we recover a threshold of $10.2\mathrm{dB}$, which is inline with similar error models Tzitrin et al. (2021).
@@
VIII.2 End-to-end simulations
As in many similar architectures Larsen et al. (2021); Fukui et al. (2018); Chamberland and Cross (2019); Bourassa et al. (2021); Tzitrin et al. (2021), the previous section abstracted the thresholds in terms of GKP effective squeezing. While this allows a comparison against different architectures, it does not give an estimate of the physical requirements for our approach. A key resource for our architecture is the squeezing level of the multi-mode Gaussian cluster state. In this section we aim to determine the threshold in terms of this resource. To this end we perform simulations of PhANTM and adaptive breeding and then use the resultant GKP states as the inputs to the logical-level simulations. We fix the PhANTM and breeding setup to use 8 PNR detectors per frequency mode with gradient beamsplitters, see Sec. IV. We chose three breeding rounds and employ a threshold for cat squeezing of $r_{lb}=0.5$ ($4.34\mathrm{dB})$, below which a momentum squeezed state from the cluster state is swapped in, as described in Sec. V. We then sweep the cluster squeezing and simulate the logical error rate for several code distances, repeating this simulation for a range of PhANTM steps.

### v3

@@ 6 Generation of Magic States
Magic states are an important ingredient of universal FTQC as they allow implementation of non-Clifford operations while themselves remaining distillable with only Clifford gates [24]. A popular gate set which generates the Clifford group is {$CNOT$, $H$, $S$}, where $CNOT$ is the controlled NOT gate; $H$ is the Hadamard gate, and $S$ is the phase gate. The non-Clifford T gate usually completes the universal gate set [24]. Provided with a source of physical magic states, one can distill a resource state that enables teleportation of non-Clifford operations into the QEC substrate at arbitrarily low error rates. Improving magic state quality by increasing the fidelity with respect to any of the |$\ket{H}$ or |$\ket{T}$ states reduces the physical resources required for a given logical error rate in the quantum computation. A promising method to produce a GKP magic state uses GKP error correction on the vacuum state [25], implying that GKP qubits and encoded Clifford operations are sufficient for universal QC [25, 71]. However, simulations show that a reasonable success rate with existing methods requires the effective squeezing of GKP states to be extraordinarily high, approaching $20$ dB (see Fig. 8), to create distillable magic states in one error correction round. In other architectures, producing magic states with practical success rates require the GKP states to have very high effective squeezing [25, 57], well above the FT threshold. In this section we provide a method to circumvent this issue and produce magic states with GKP resources that have effective squeezing levels near the FT threshold.
@@
The distillation infidelity threshold for H-type states is approximately 14.7%. The plots in Fig. 8 show a deterministic simulation of CV-QEC on input vacuum or cat states with post-selection on quadrature measurement outcomes of zero. Note that this post-selection, done to ease simulations, will slightly underestimate the output squeezing of the bred GKP states [54]. We see that when performing CV-QEC on the vacuum mode (purple data points) one obtains infidelities above the distillation threshold (of 14.7%) when using GKP states with squeezing of 13dB or less as the resource state for CV-QEC. On the other hand, with effective squeezing greater than 12dB and cat states with amplitude of $\sim 5.1$, magic states with infidelities below the distillation threshold can be obtained.
@@
In practice, the cat states produced by PhANTM have random amplitudes, and they must be squeezed in order to obtain the correct spacing for GKP breeding before magic state generation. This means that our GKP states have a random squeezing distribution. Further, as Fig. 8 shows, the optimal input cat amplitude for the error-corrected mode is between 4.9 and 5.5. Because our cats have random amplitudes and squeezing levels, only a few will hit this required target, although PhANTM can be engineered to emit cat states with mean squeezing levels and amplitudes centered on the requirements. Monte Carlo simulations were used to determine the success rate of producing a magic state when random cats from PhANTM were used to provide the resource states, both for GKP breeding and the cat state upon which CV QEC was performed. The simulations use Qutip [69] for density matrix calculations and Strawberry Fields [70] for homodyne detection simulation. Fig. 10 shows histograms of the infidelities obtained from each magic state generation attempt, where the fidelity is measured with respect to an ideal $H$-type magic state. The cat states were produced using Monte Carlo PhANTM simulations with a cluster state squeezing of 13.5 dB, 20 PhANTM steps, and 8 photon-subtraction attempts made before homodyne detection and teleportation to the next PhANTM time step. When using vacuum states as the error-corrected mode, we see a success rate of 0.003 after 1000 iterations (brown counts in the histogram). On the other hand, when using cat states as the error-corrected mode, we see a success rate of 0.048 after 1000 magic state generation attempts, an improvement of more than factor of 10, which reduces resources required to produce a magic state by a factor 10 relative to using vacuum alone. This simulation specifically targeted $H$-type states, which enable teleportation of $T$ gates within the architecture. These success rates mean that magic factories will not require the majority of qubit resources in surface code substrates using lattice surgery [72].
@@
7 Encoding Logical Qubits in a Surface Code
GKP encoding can protect against small displacement errors, but larger displacements ($>\frac{\pi}{2}$) will result in an error. Depending on the quadrature, this can be either a Pauli X error ($q$ quadrature) or a Pauli Z error ($p$ quadrature). These logical qubit errors require additional qubit-level error-correction. Many error correcting codes have been studied in the context of GKP qubits [26, 73, 74, 75]. We will focus on topological error correction based on a family of codes known as surface codes, which have been shown to have high thresholds ($\approx 1\%$ under circuit level noise [76] or $\approx 10\mathrm{dB}$ with GKP states [26]) and well-studied logical gates [77].
As mentioned earlier, photonic CVQC is well-suited to MBQC. Circuit model codes known as Calderbank-Steane-Short(CSS) codes, where X and Z errors are detected independently, can be translated into a cluster state for fault-tolerant MBQC via a process known as foliation [78], which loosely maps the data qubits in the circuit model code to linear cluster states. Surface codes are examples of CSS codes, and the resulting foliated 3D cluster state is often called the Raussendorf-Harrington-Goyal (RHG) lattice after the authors who first introduced it [79]. The unit cell of this cluster state is shown in Fig. 11. The product of $X$-basis measurements (equivalently $p$-basis homodyne measurements for GKP qubits) on all face qubits is a stabilizer of the cluster state. In the error free case, measuring stabilizers will result in +1 eigenvalue and errors will result in $-1$ measurement results for any stabilizer that they anti-commute with. A second unit cell can be defined by shifting the base unit cell by 1/2 in all directions. Here, the edge qubits of the initial unit cell become face qubits of the second unit cell. These two lattices are known as primal and dual lattices, respectively. By measuring the stabilizers of both lattices, errors on all qubits can be detected. Logical operators in foliated cluster states are formed by correlation surfaces that map logical operators from the underlying code from one end of the cluster state to the other. To implement logical X and Z operations, correlation surfaces of the RHG lattice are given by 2D sheets of qubits which span the lattice and intersect primal and dual unit cells, respectively. The code distance is the shortest chain of errors that can result in a logical error, which is equivalent to the shortest error chain that can span the RHG lattice. The threshold theorem states that provided the physical error rate of the underlying qubits is below some threshold, we can arbitrarily suppress the logical error rate by increasing the code distance [80].
@@ 7 Encoding Logical Qubits in a Surface Code
The popular Clifford + T gate set maps naturally to the surface code. Clifford gates and teleportation based T gates can be implemented by lattice surgery [77]. This approach defines logical qubits as patches, which are subsections of the overall surface code lattice, with logical operators identified on the edges. In MBQC, this corresponds to volumes of the RHG lattice with correlation surfaces on the boundaries. Patches can be merged by measuring stabilizer checks between edges of the patches, thus performing a joint logical measurement of the two patches. Clifford gates and the T gate can be implemented with joint measurements including ancilla qubits or magic states. In fact, in Ref. [77], it was shown how all Clifford gates can be commuted through the T gates and measurements, resulting in circuits that just consist of mulit-qubit joint measurements including a magic state. Lattice surgery operations can also be mapped to changing measurement bases (homodyne angles) in fault-tolerant MBQC [61].
This gate set is well studied and optimized, however our architecture can be adapted to other gate sets [77, 81] and other error correcting codes [82]. In this paper we aim to quantify Gaussian squeezing resources needed to demonstrate FT thresholds with the RHG lattice serving as the example code. We leave optimization of the error correcting code and gate set for future work.
7.1 Macronode GKP lattice construction
@@
We can construct the lattice directly, where the nodes are GKP $\ket{+}$ states and the edges are CV $C_{Z}$ gates, which can be implemented deterministically with linear optics and squeezing. This approach has two main draw backs. Firstly, while it has been demonstrated previously [83], in-line squeezing is generally difficult to realize experimentally. Second, ideal $C_{Z}$ gates are not physical, requiring infinite squeezing. Approximating the gate with finite squeezing will add noise to the GKP qubits. Alternatively, any lattice can be constructed with passive interferometers starting with GKP Bell pairs, where the number of physical modes per macronode is the valency of the underlying canonical graph state and the GKP Bell pairs are distributed to match the entanglement in the canonical graph state. Relating measurements on the physical modes with effective measurements on the underlying lattice requires a list of corrections known as a dictionary protocol [58], which we describe in Appendix 11.6. Note, that in the following we use the terms GKP state and sensor state interchangeably, always meaning either a sensor state or one half of a GKP Bell pair formed by interfering sensor states.
@@
We wish to design a modular construction of the required lattice. To this end, we identify a unit cell consisting of 24 sensor states or 12 GKP Bell pairs that can be tiled to construct a macronode RHG lattice of arbitrary size. From this unit cell, four GKP nodes are sent to neighboring unit cells, and correspondingly four are received. Two of the modes are delayed to form temporal links. This is illustrated in Fig. 11. As the primal and dual layers are created in a single time-step, we avoid the requirement for adaptive optics that have often been used to account for the different entanglement structure in primal and dual layers [26, 28]. To teleport non-Clifford operations onto logical qubits within the surface code, physical magic states are injected into the RHG lattice [72, 61]. The physical magic state is interfered with a vacuum state at a beamsplitter and is used in place of one GKP Bell pair in the RHG macronode.
We note here that due to the inherent randomness in the measurements, there are several feed forward corrections that need to be accounted for. Some of them, such as the corrections due to the macronode construction and the random outcomes on the effective Pauli measurements on the RHG lattice, do not need to be physically implemented and can be tracked in software. While certain operations do require physically implementing the feedforward, such as changing measurement bases due to magic state injection, changing the measurement basis corresponds to changing the homodyne phase for GKP based architectures. In this way, the required adaptivity at the QEC level does not require any fast modulators on the qubit modes.
8 FTQC Threshold Simulation
@@
Several CV FTQC architectures based on GKP qubits and the surface code exist in the literature and have been characterised by their GKP effective squeezing threshold. Refs. [84] and [26] used constructions based on $C_{Z}$ gates and found thresholds of $10$ dB and $10.5$ dB respectively, however these ignored the additional noise from finite squeezing in physical gates. In Ref. [28] found a threshold of $12.7$ dB while accounting for finite squeezing. Ref. [27] proposed a macronode-based architecture which did away with the need for $C_{Z}$ gates and found a threshold of $10.1$ dB. Refs. [26, 28, 27] all require fast reconfigurable optics either in the computation or in the preparation of GKP states with a sufficiently high success probability.
@@
In the following section, we analyze the fault-tolerant performance of our architecture. In order to allow comparison with other architectures, we first use a phenomenological error model used widely in the literature [84, 85, 28], which models realistic GKP states as ideal GKP states acted on by a Gaussian displacement channel. Formally, this is represented as [27]
@@
Which has the effect of replacing the delta function teeth of an ideal GKP state with Gaussian distributions with a standard deviation, $\Delta$,where $\Delta$ describes the same broadening of the GKP teeth as $\Delta_{D}$ described in Sec. 2 [26]. We can link this to the GKP effective squeezing in dB as $S_{\mathrm{dB}}=-\log_{10}\left(\Delta^{2}\right)$ [26]. Despite its simplicity, this error model can describe many common noise sources such as finite squeezing and photon loss [27, 26, 28, 86]. We then move on to quantify the fault-tolerance performance in terms of the input Gaussian resource for PhANTM - the squeezing level of the multi-mode cluster state. The input for these simulations are distributions of GKP squeezings obtained from simulations of the proposed GKP generation pipeline.
@@
In both cases, to determine the logical error rate we track how the GKP effective squeezing propagates through the optical circuit used to generate the macronode lattice and use the propagated noise variances to simulate noisy homodyne measurements. We apply the corresponding post-processing to reduce the physical macronode measurements to the effective noisy measurements on the canonical lattice, as shown in Appendix 11.6. With these noisy measurements we determine the probability of a qubit level error on each node of the canonical lattice, see Appendix 11.7 for more detail. These probabilities are used to sample a set of errors and generate the corresponding syndrome of the RHG lattice. This syndrome is then decoded using the pyMatching implementation of the minimum weight perfect matching decoder [87]. This simulation method is described in more detail in Appendix 11.8 and is similar to the method used in Ref [26], adapted to the macronode approach in Ref. [27] and for our specific GKP generation method.
8.1 Phenomenological error model
@@
Here, we present results which tailor the phenomenological noise model to the PhANTM and adaptive breeding approach to generate GKP states. As described in Sec. 5, generating GKP states from PHANTM and breeding results in GKP states that are inherently asymmetric due to the different physical processes that determine the squeezing in the GKP quadratures. Note that this is equivalent to grid states with asymmetric spacing in the two quadratures [73] which has also been studied in [82]. This feature was incorporated into our simulation method by defining $p$ and $q$ noise variances separately for each GKP state. While we can also set each GKP state independently, for simplicity we assume that that the $p$ and $q$ noise are identical for each GKP state. To determine the 2D threshold curve for unbalanced quadrature noise we employ the method used in Ref. [13]. By expressing the $p$ and $q$ noise in terms of a dummy variable, $\Delta_{p}=c_{p}\times x$, $\Delta_{q}=c_{q}\times x$, we fix the ratio of $p$ and $q$ noise via $c_{p}/c_{q}$ and use $x$ to find the threshold for a given ratio.
@@
Fig. 12 shows the resultant surface, where we have shaded the correctable region. The surface is symmetric with marginal squeezing thresholds of $7.2\mathrm{dB}$. With balanced squeezing in each quadrature we recover a threshold of $10.2\mathrm{dB}$, which is inline with similar error models [27].
@@
8.2 End-to-end simulations
As in many similar architectures [28, 84, 88, 26, 27], the previous section abstracted the thresholds in terms of GKP effective squeezing. While this allows a comparison against different architectures, it does not give an estimate of the physical requirements for our approach. A key resource for our architecture is the squeezing level of the multi-mode Gaussian cluster state. In this section we aim to determine the threshold in terms of this resource. To this end we perform simulations of PhANTM and adaptive breeding and then use the resultant GKP states as the inputs to the logical-level simulations. We fix the PhANTM and breeding setup to use eight PNR detectors per frequency mode with gradient beamsplitters, see Sec. 4. We chose three breeding rounds and employ a threshold for cat squeezing of $r_{lb}=0.5$ ($4.34\mathrm{dB})$, below which a momentum squeezed state from the cluster state is swapped in, as described in Sec. 5. We then sweep the cluster squeezing and simulate the logical error rate for several code distances, repeating this simulation for a range of PhANTM steps.